OpenStudy (jayden):

Solve 4(x + 1) + 8 = 24. (1 point) 3.75 -2 5 3 please help

OpenStudy (3mar):

May I help?

OpenStudy (jayden):

yes please

alones (alones):

Sart from simplify both sides

OpenStudy (3mar):

\[4(x+1)+8=24\]

OpenStudy (3mar):

then we multiply 4 by the parentheses: \[4x+4+8=24\]

OpenStudy (3mar):

add 4 to 8: \[4x-12=24\]

OpenStudy (3mar):

sorry

OpenStudy (jayden):

so x=3 thanks

OpenStudy (3mar):

\[4x+12=24\]

OpenStudy (3mar):

Yes, correct! x=3